### v2

@@ 3.2 Results
This section provides a detailed analysis of various data management aspects and answers the two RQs. First, we define four data management aspects, as identified in our SLR in Section 3.2.1. Structured based on these aspects, we then discuss the challenges (RQ1) and solutions (RQ2) associated with each aspect in detail (see Sections 3.2.33.2.6). Note that we discuss the challenges and solutions together, as this will provide readers with a better context of each challenge and the related solution as discussed in the associated study.
@@ 3.2.1 Data Management Aspects
Data Collection is an activity that involves gathering and distributing information. This information is subsequently utilized for analysis and decision-making, as mentioned in the studies discussed data collection aspect in Table 1.
@@ 3.2.2 Overlapping Data Management Aspects
Figure 3 illustrates the connections between various data management aspects in agile software development, revealing that addressing a single challenge can lead to the resolution of other connected challenges. This insight underscores the importance of implementing integrated and comprehensive solutions to enhance the effectiveness and efficiency of agile projects.
@@
Data Integration and Data Quality: The studies depicted in Figure 3 discussed challenges and solutions related to data integration and data quality, showing that addressing data integration challenges, as outlined in Section 3.2.3, inherently resolves data quality issues, which are discussed in Section 3.2.5. For example, the solution proposed by Spengler et al. spengler2020enabling (cloud-based platform with a generic agile data-loading pipeline) integrates data from distributed and heterogeneous sources. Additionally, this pipeline is designed to automatically detect the syntax and format of input data, handle different encodings, and manage missing and duplicate data. Another example, the solution proposed by Rosenkranz et al. rosenkranz2017supporting addresses the challenges in both data integration (harmonizing data from different sources) and data quality (ensuring accuracy and consistency), highlighting the interconnected nature of these two aspects in effective data management.
@@ 3.2.3 Data Integration Challenges and Solutions
Table 2: Summary of Data Integration Challenges and Solutions
Data Integration Challenge Data Integration Solution Solution Status Ref.
@@
Translating the metadata of the datasets into different formats Implemented schuttler2021journey
@@
Solution referenced in Section 3.2.4 Partially implemented upender2005staying
Data Harmonization and Interoperability: Abdallah and Fan abdallah2022towards discussed the challenge of harmonization and interoperability between heterogeneous Maintenance Repair and Overhaul (MRO) information systems, which is a critical issue in integrating maintenance records over the lifespan of an aircraft. Spengler et al. spengler2020enabling discussed the challenge of integrating data from distributed and heterogeneous sources at the technical, structural, and semantic levels. Vøgt et al. vogt2023implementing addressed incorporating new climate data into local government processes. Schüttler et al. schuttler2021journey focused on coordinating the definition and integration of datasets to ensure interoperability and avoid the development of parallel structures, necessitating alignment with other research infrastructures like the European Biobanking and BioMolecular Resources Research Infrastructure - European Research Infrastructure Consortium (BBMRI-ERIC) and the German Medical Informatics Initiative (MII), which is a national initiative aimed at improving medical research and healthcare by facilitating the sharing and integration of data across various medical institutions.
@@
Semantic Heterogeneity: refers to the challenge of integrating data with divergent interpretations and meanings. Rosenkranz et al. rosenkranz2017supporting highlighted the problem of semantic heterogeneity during the design of data integration requirements in the development of data warehouses (DW). Dos Santos Júnior et al. dos2021ontology discussed the difficulty in accessing, integrating, analyzing, and viewing data across heterogeneous applications with varying semantics. Abdallah and Fan abdallah2022towards discussed semantic heterogeneity in heterogeneous MRO systems. Barcellos and Perini barcellos2020towards discussed the difficulty of integrating data from different agents and tools due to semantic heterogeneity. Rix et al. rix2016agile discussed the difficulty of integrating and making sense of diverse data sources with varying formats and meanings within the high-pressure die casting (HPDC) manufacturing process.
@@
To resolve the semantic heterogeneity challenges, Rosenkranz et al. rosenkranz2017supporting proposed a communication-centric approach that combines agile software development practices with communication theory. This approach includes three core artifacts: a detailed template for data field specifications, a procedure model for iterative development, and the use of agile methods to facilitate continuous stakeholder communication. These components work together to enhance stakeholder comprehension and ensure consistent data usage across the organization. The solution was implemented and evaluated. Dos Santos Júnior et al. dos2021ontology proposed developing an ontology-based approach known as Immigrant as a solution. By mapping the diverse semantics of various applications onto a common ontology, Immigrant seeks to resolve semantic conflicts and provide a cohesive and integrated view of the data. This enables the presentation of meaningful information through dashboards, facilitating data-driven decision-making in agile software organizations. The approach is currently in development, with plans for a proof of concept to validate its effectiveness. In addition to the data harmonization and interoperability challenge addressed by Abdallah and Fan abdallah2022towards suggested the development of ontologies discussed above to resolve data harmonization, interoperability, and semantic heterogeneity in heterogeneous MRO systems.
@@
Barcellos and Perini barcellos2020towards proposed using ontologies to establish a common conceptualization across different tools and data sources. This approach helps to reduce semantic conflicts and enables correct data integration by acting as an interlingua to map the concepts used by various agents and tools, thus facilitating understanding and communication. They developed and used domain ontologies as reference models to integrate tools that support Continuous Software Engineering (CSE) processes. They started by developing and integrating a Scrum Reference Ontology with ontologies from the Software Engineering Ontology Network (SEON), covering requirements and project management aspects. This ontology was then applied as a basis to integrate tools such as Clockify and Azure DevOps used by development teams in the software unit of a Brazilian government agency. As a result, data from different tools were integrated and displayed in dashboards, providing useful information for managers to make decisions. Rix et al. rix2016agile proposed a solution that implemented a modular and agile information processing framework. This framework utilizes annotation services, which add metadata to data sources, and a Master Data Management (MDM) repository, a centralized database that stores and manages master data. It enables real-time data streaming for data mining analyses and visualization of results. The framework ensures that data from various sources are annotated with metadata, facilitating pervasive traceability and simplifying information aggregation. Additionally, using dynamic Object-Relational Mapping (ORM) frameworks, such as Hibernate, in conjunction with open interfaces like OPC-UA (Open Platform Communications Unified Architecture), addresses semantic heterogeneity. This is accomplished by enabling the labeling and exchange of information in a unified manner. The solution was implemented and evaluated at the Audi testing foundry.
@@
Data Transformation and Extraction: Data transformation and extraction involve converting and transferring data from various sources into a usable format for analytics and decision-making. Hofer et al. hofer2020new the challenge of extracting and refining data from semi-structured, mixed-quality crowd-sourced data such as Wikipedia and Wikidata. Dursun et al. dursun2014workflow emphasized the need for a unified view of data from various sources and to integrate data from heterogeneous stores into a single, coherent data store for analytics and visualization. Kannan et al. kannan2017rapid discussed the importance of an efficient ETL process for extracting data from Electronic Health Records (EHR) into an Enterprise Data Warehouse (EDW). Dharmapal et al. dharmapal2016big highlighted the complexity of integrating diverse data types without complex, time-consuming IT engineering efforts and suggested that modern analytic processing tools should transition away from traditional, retrospective BI tools and platforms towards more progressive analytic platforms.
@@
Managing Data Integration: Vestues et al. vestues2022agile discussed the difficulty of managing data across various silos within the organization and supporting analytical solutions that require cross-organizational insights. Little et al. little2004adaptive mentioned that a key business value of Landmark Graphics (a leading supplier of software and services) application suite arises from the integration of these products through a standard data model with over 800 tables, 12 entities, and 90 data types. This complexity presents a challenge in managing and integrating data across different products. Chen et al. chen2016agile discussed the challenge of effectively managing internal and external data integration in big data projects, including data from existing Enterprise Data Warehouses (EDW) and new NoSQL systems.
@@
To address the challenges related to managing data integration, Vestues et al. vestues2022agile proposed the adoption of a “data mesh” approach. This approach comprises four core principles: domain-oriented decentralized data ownership and architecture, data as a product, self-serve data infrastructure, and federated computational governance. The data mesh model advocates for a shift from centralized data management to a distributed model where domain teams have increased ownership of the data produced by their applications, aiming to improve coordination and enable more agile and automated approaches to data analytics. Vestues et al. vestues2022agile focused on reporting findings from a case study of a public sector organization in Norway that has begun the transition from centralized to distributed data management. Little et al. little2004adaptive presented that Landmark Graphics adopted the Microsoft Solutions Framework (MSF), a milestone-based iterative development framework, to standardize their development process across different product teams. To manage the integration of diverse data types effectively of integrated products and ensure consistency in the integration efforts, which was crucial given the complexity of their common data model. Chen et al. chen2016agile adopted an architecture-centric approach, which strongly emphasizes designing and developing a system’s architecture as a crucial element of the development process. This approach entails the creation of a well-defined architecture, which acts as a blueprint for the system’s development. It guides the selection of technologies, allocation of resources, and risk management throughout the development life cycle. This approach complements agile’s modular design, a software development approach that emphasizes breaking down a project into smaller, more manageable modules or components to effectively manage the integration of diverse data types. The approach includes using data lakes and cloud storage to handle various data types, supporting agile’s requirement for flexible data structures and continuous delivery. The new methodology, AABA (Architecture-centric Agile Big Data Analytics), was implemented and validated through multiple case studies encompassing 10 big data analytics projects.
@@
Diverse and Decentralized Data sources: Upender upender2005staying discussed the complexity involved in integrating patient data and transitioning to a centralized data management system within a diverse and decentralized user environment. Harper and Dagnino harper2014agile highlighted the need to merge and clean data sets from different sources to enable advanced data analytics. The challenge is initiating and sustaining processes while ensuring data exchange and requests between components. The proposed solution is to distribute work requests across computing resources. Chhillar and Sharma chhillar2019act discussed the challenge of consolidating reports from various sources, such as development, testing, and bug-tracking tools, into a unified data analytics tool. This challenge pertains to seamlessly integrating diverse data sources and formats to generate real-time metrics and reports for software quality analysis.
@@ 3.2.4 Data Collection Challenges and Solutions
Table 3: Summary of Data Collection Challenges and Solutions
Data Collection Challenge Data Collection Solution Solution Status Ref.
@@
Capturing Diverse Data User-centered design strategies for data collection Implemented pater2018advancing
Continuous Integration servers and other software development tools Proposed matthies2019towards
Analytics-Driven Testing (ADT) process Proposed batarseh2018predicting
@@
Data Collection Method Challenges Automated Continuous Testing TestBot and ACQ Metrics dashboard Partially Implemented chhillar2019act
Automation of data collection and visualization toolchain Proposed lehtonen2017visualizations
Human-centred Agile Software Engineering (HASE) platform Proposed lin2018towards
Automated methods for data collection and visualization Proposed svensson2019unfulfilled
Centralized data management system with Scrum and XP practices Proposed upender2005staying
@@
Data Sharing and Collaboration Diagnostic model for understanding data collection and sharing practices Proposed fabijan2016lack
Data-Driven Systems Engineering (DDSE) methodologies Proposed kaur2020dialogue
Legal advisor for data sensitivity assessment Proposed barbala2023data
Informative Data collection challenges Development of ontology-based approach ’Immigrant’ Proposed dos2021ontology
@@
Integration of development project data into Retrospective agendas Proposed matthies2020playing
Q-Rapids tool for acquiring valuable information Implemented martinez2019continuously
Agile methodology and Data Analyst collaboration for data collection Proposed dharmapal2016big
@@
Comprehensive Data Collection Creation of a generalized dataset and standardized data processing Proposed das2015towards
Big data stack with scalable storage technologies Proposed huang2021leveraging
Project Management Information System for automated data collection Proposed fagarasan2023integrating
Automated data collection mechanisms from software in use Proposed olsson2018challenges
@@
Capturing Diverse Data: Pater et al. pater2018advancing focused on capturing data for personalized healthcare and human services while addressing barriers to widespread adoption and standardization of information systems within healthcare. Matthies and Hesse matthies2019towards discussed the need for effective data collection and analysis from diverse sources, including software project data, customer feedback, and operational metrics. Furthermore, Batarseh and Gonzalez batarseh2018predicting highlighted the issue of data is often not collected in a structured and organized manner, making it difficult to analyze and use for decision-making.
@@
To address capturing diverse data challenges, Pater et al. pater2018advancing proposed a solution that involves leveraging a variety of user-centered design strategies, such as participatory design, empathic design, and design thinking, to guide the data collection methods. This approach maintains the user at the center of the process and continuously incorporates user-driven input into each phase of the agile development cycle. The solution was implemented to modernize the Army Community Service’s information technology infrastructure. Matthies and Hesse matthies2019towards highlighted tools such as Continuous Integration servers, static analysis tools, and other software development tools that provide valuable data points on the current status and health of the developed software project. They highlighted that combining data analysis and interpretation by agile teams can enable better-informed decisions in business and software development processes. Batarseh and Gonzalez batarseh2018predicting proposed the Analytics-Driven Testing (ADT) process. ADT involves a structured data collection and analysis approach within the agile software development lifecycle. During each sprint, data are systematically collected using the ADT tool. This includes recording all software failures, measuring the Mean Time Between Failures (MTBF), and collecting other relevant data such as software module, sprint number, and run time.
@@
Data Collection Method Challenges: Chhillar and Sharma chhillar2019act discussed the difficulty in generating real-time software quality metrics and the need for actual data collection processes. In contrast, Lehtonen et al. lehtonen2017visualizations discussed the manual nature of data collection from the issue management system, which can be time-consuming and susceptible to errors. Upender upender2005staying discussed the challenge involves creating a system that allows for the electronic capture of clinical research data, replacing the existing rudimentary tools such as Excel and Access that various independent research groups use. Lin et al. lin2018towards discussed the difficulty of collecting data on software engineering activities, such as task allocation, collaboration, and mood stability, without disrupting the normal Agile Software Development (ASD) process. Practitioners surveyed by Svensson et al. svensson2019unfulfilled expressed concerns regarding data availability, excessive quantity, and ambiguity in terms of usability, relevance, and its connection to decision-making.

### v4

@@ 3.2 Results
This section provides a detailed analysis of various data management aspects and answers the two RQs. First, we define four data management aspects, as identified in our SLR in Section 3.2.1. Structured based on these aspects, we then discuss the challenges (RQ1) and solutions (RQ2) associated with each aspect in detail (see Sections 3.2.43.2.7). Note that we discuss the challenges and solutions together, as this will provide readers with a better context of each challenge and the related solution as discussed in the associated study. We provided a definition for each of those challenges and solutions, which emerged from analyzing and synthesizing the studies included in the SLR.
@@ 3.2.1 Data Management Aspects
Data Collection is an activity that involves gathering information. This information is subsequently utilized for analysis and decision-making, as mentioned in the studies discussed data collection aspect in Table 1.
@@ 3.2.2 Overlapping Data Management Aspects
Figure 3 illustrates the connections between various data management aspects in agile software development, revealing that addressing a single challenge can resolve other connected challenges. This insight underscores the importance of implementing integrated and comprehensive solutions to enhance the effectiveness and efficiency of agile projects.
@@
Data Integration and Data Quality: The studies depicted in Figure 3 discussed challenges and solutions related to data integration and data quality, showing that addressing data integration challenges (Section 3.2.4), inherently resolves data quality issues (Section 3.2.6). For example, the solution proposed by Spengler et al. spengler2020enabling (cloud-based platform with a generic agile data-loading pipeline) integrates data from distributed and heterogeneous sources. Additionally, this pipeline is designed to automatically detect the syntax and format of input data, handle different encodings, and manage missing and duplicate data. Another example, the solution proposed by Rosenkranz et al. rosenkranz2017supporting addresses the challenges in both data integration (harmonizing data from different sources) and data quality (ensuring accuracy and consistency), highlighting the interconnected nature of these two aspects in effective data management.
3.2.3 Data Management Challenges Prioritization
We prioritized the data management challenges based on their prevalence in reviewed studies and their potential impact on agile workflows (see Table 2).
Prevalence was determined by counting the number of studies discussing each challenge. This approach aligns with systematic review guidelines kitchenham2007guidelines , which prioritize frequently discussed issues as indicators of their importance. For example, as shown in Table 2, data harmonization and interoperability challenges appeared in 4 studies, highlighting its recurring relevance across domains. Its frequent mention underscores its important role in enabling consistency and seamless integration of diverse systems, a common need across agile projects. Challenges like this, which were discussed extensively, demonstrate their widespread recognition and applicability in practical agile environments.
The potential impact was assessed based on how each challenge influences agile practices, such as sprint planning, iterative development, and cross-team collaboration. We categorized impact as follows:
  • High: Challenges directly disrupting these practices, such as delaying decision-making or impeding team alignment, were rated high impact.
  • Medium: In contrast, those with less severe effects were rated as medium impact. For example, ensuring data accuracy and consistency challenge was identified as having a high impact in 5 studies , as inaccurate or inconsistent data can mislead sprint planning and decision-making, causing project delays and inefficiencies. In contrast, while important, Semantic Heterogeneity was rated as medium impact because its effects can often be mitigated with mechanisms such as the use of ontologies.
Note that all of the identified challenges were found to have at least medium significance due to their impact on agile workflows. No challenges were categorized as ”low impact,” reflecting the focus on issues with meaningful implications for agile practices. This highlights the importance of data management in agile environments, where even moderate challenges can significantly affect project outcomes.
Table 2: Prioritized Data Management Challenges
Data Management Challenges
# Studies
Potential Impact on Agile Workflows
Data Integration
Data Harmonization and Interoperability 4 High – Essential for cross-team workflows
Semantic Heterogeneity 4 Medium – Critical for accurate decision-making
Data Transformation and Extraction
4 High – Impacts real-time data usability
Managing Data Integration 3 Medium – Addresses cross-functional coordination
Diverse and Decentralized Data Sources 3 High – Influences accessibility and consistency
Data Collection
Capturing Diverse Data 3 High – Impacts completeness of agile decisions
Comprehensive Data Collection 3 Medium – Enables broader data applicability
Data Sharing and Collaboration 2 Medium – Improves team alignment
Informative Data Collection 2 Medium – Supports specific project needs
Data Quality
Ensuring Data Accuracy and Consistency 5 High – Critical for decision reliability
Missing Quality Data 3 High – Affects agile sprint outputs
Inadequate Data Quality Management 3 Medium – Increases the risk of errors
Data Quality Standardization 2 Medium – Facilitates uniform data processes
Data Analysis
Analyzing Large and Complex Data 5 High – Influences scalability and insights
Analyzing Semantic Heterogeneity Data 3 Medium – Ensures semantic alignment
Efficient Data Analysis and Visualization 3 High – Impacts decision-making speed
Real-Time Data Analytics and Decision Making 2 High – Supports adaptive project actions
Selection of Appropriate Analytical Techniques 2 Medium – Improves analysis relevance
@@ 3.2.4 Data Integration Challenges and Solutions
Table 3: Summary of Data Integration Challenges and Solutions identified in the SLR.
Challenge Solution Solution Status Ref.
@@
Translating the datasets’ metadata into different formats Implemented schuttler2021journey
@@
Ref in Sec. 3.2.5 Partially implemented upender2005staying
Data Harmonization and Interoperability: Data harmonization ensures consistency across diverse sources, while interoperability enables seamless data exchange between systems. In agile development, these processes are crucial for integrating data across teams and tools amid frequent iterations and dynamic requirements. Without them, teams face inefficiencies like redundant efforts, fragmented datasets, and poor data alignment, hindering decision-making and project agility. Abdallah and Fan abdallah2022towards discussed the challenge of harmonization and interoperability between heterogeneous Maintenance Repair and Overhaul (MRO) information systems, which is a critical issue in integrating maintenance records over the lifespan of an aircraft. Spengler et al. spengler2020enabling discussed the challenge of integrating data from distributed and heterogeneous sources at the technical, structural, and semantic levels. Vøgt et al. vogt2023implementing addressed incorporating new climate data into local government processes. Schüttler et al. schuttler2021journey focused on coordinating the definition and integration of datasets to ensure interoperability and avoid the development of parallel structures, necessitating alignment with other research infrastructures like the European Biobanking and BioMolecular Resources Research Infrastructure - European Research Infrastructure Consortium (BBMRI-ERIC) and the German Medical Informatics Initiative (MII), which is a national initiative aimed at improving medical research and healthcare by facilitating the sharing and integration of data across various medical institutions.
@@
Semantic Heterogeneity: This refers to the challenge of integrating data with differing meanings or interpretations. In agile development, where teams often work with diverse tools and datasets, conflicting data semantics can lead to misalignment in decision-making and hinder collaboration. Resolving these inconsistencies is essential to ensure accurate data integration and good understanding within teams. Barcellos barcellos2020towards discussed the difficulty of integrating data from different agents and tools due to semantic heterogeneity. Rix et al. rix2016agile discussed the difficulty of integrating and making sense of diverse data sources with varying formats and meanings within the high-pressure die casting (HPDC) manufacturing process.
@@
To resolve the semantic heterogeneity challenges, Rosenkranz et al. rosenkranz2017supporting proposed a communication-centric approach that combines agile software development practices with communication theory. This approach includes three core artifacts: a detailed template for data field specifications, a procedure model for iterative development, and agile methods to facilitate continuous stakeholder communication. These components work together to enhance stakeholder comprehension and ensure consistent data usage across the organization. The solution was implemented and evaluated. Dos Santos Júnior et al. dos2021ontology proposed developing an ontology-based approach known as Immigrant as a solution. By mapping the diverse semantics of various applications onto a common ontology, Immigrant seeks to resolve semantic conflicts and provide a cohesive and integrated view of the data. This enables the presentation of meaningful information through dashboards, facilitating data-driven decision-making in agile software organizations. The approach is currently in development, with plans for a proof of concept to validate its effectiveness. In addition to the data harmonization and interoperability challenge addressed by Abdallah and Fan abdallah2022towards suggested the development of ontologies discussed above to resolve data harmonization, interoperability, and semantic heterogeneity in heterogeneous MRO systems.
@@
Barcellos barcellos2020towards proposed using ontologies to establish a common conceptualization across different tools and data sources. This approach helps to reduce semantic conflicts and enables correct data integration by providing a shared conceptual framework to align and map the concepts used by various agents and tools, thus facilitating understanding and communication. They developed and used domain ontologies as reference models to integrate tools that support Continuous Software Engineering (CSE) processes. They started by developing and integrating a Scrum Reference Ontology with ontologies from the Software Engineering Ontology Network (SEON), covering requirements and project management aspects. This ontology was then applied as a basis to integrate tools such as Clockify and Azure DevOps used by development teams in the software unit of a Brazilian government agency. As a result, data from different tools were integrated and displayed in dashboards, providing useful information for managers to make decisions. Rix et al. rix2016agile proposed a solution that implemented a modular and agile information processing framework. This framework utilizes annotation services, which add metadata to data sources, and a Master Data Management (MDM) repository, a centralized database that stores and manages master data. It enables real-time data streaming for data mining analyses and visualization of results. The framework ensures that data from various sources are annotated with metadata, facilitating pervasive traceability and simplifying information aggregation. Additionally, using dynamic Object-Relational Mapping (ORM) frameworks, such as Hibernate, in conjunction with open interfaces like OPC-UA (Open Platform Communications Unified Architecture), addresses semantic heterogeneity. This is accomplished by enabling the labeling and exchange of information in a unified manner. The solution was implemented and evaluated at the Audi testing foundry.
@@
Data Transformation and Extraction: Data transformation involves converting data from diverse formats and sources into a usable form for analysis and decision-making. Agile teams often need to handle real-time data, requiring efficient processes to clean, format, and integrate information. Ineffective transformation and extraction can delay insights, disrupt workflows, and impede the agile principle of continuous delivery. Hofer et al. hofer2020new the challenge of extracting and refining data from semi-structured, mixed-quality crowd-sourced data such as Wikipedia and Wikidata. Dursun et al. dursun2014workflow emphasized the need for a unified view of data from various sources and to integrate data from heterogeneous stores into a single, coherent data store for analytics and visualization. Kannan et al. kannan2017rapid discussed the importance of an efficient ETL process for extracting data from Electronic Health Records (EHR) into an Enterprise Data Warehouse (EDW). Dharmapal et al. dharmapal2016big highlighted the complexity of integrating diverse data types without complex, time-consuming IT engineering efforts and suggested that modern analytic processing tools should transition away from traditional, retrospective BI tools and platforms towards more progressive analytic platforms.
@@
Managing Data Integration: This requires combining and aligning data from multiple systems into a unified dataset. Frequent updates and evolving requirements in agile environments add complexity to this process. Without an effective integration process in place, teams may struggle to generate cross-organizational insights, leading to inefficiencies and limited data-driven decision-making. Vestues et al. vestues2022agile discussed the difficulty of managing data across various silos within the organization and supporting analytical solutions that require cross-organizational insights. Little et al. little2004adaptive mentioned that a key business value of Landmark Graphics (a leading supplier of software and services) application suite arises from the integration of these products through a standard data model with over 800 tables, 12 entities, and 90 data types. This complexity presents a challenge in managing and integrating data across different products. Chen et al. chen2016agile discussed the challenge of effectively managing internal and external data integration in big data projects, including data from existing Enterprise Data Warehouses (EDW) and new NoSQL systems.
@@
To address the challenges related to managing data integration, Vestues et al. vestues2022agile proposed the adoption of a “data mesh” approach. This approach comprises four core principles: domain-oriented decentralized data ownership and architecture, data as a product, self-serve data infrastructure, and federated computational governance. The data mesh model advocates for a shift from centralized data management to a distributed model where domain teams have increased ownership of the data produced by their applications, aiming to improve coordination and enable more agile and automated approaches to data analytics. Vestues et al. vestues2022agile focused on reporting findings from a case study of a public sector organization in Norway that has begun the transition from centralized to distributed data management. Little et al. little2004adaptive presented that Landmark Graphics adopted the Microsoft Solutions Framework (MSF), a milestone-based iterative development framework, to standardize their development process across different product teams. To manage the integration of diverse data types effectively of integrated products and ensure consistency in the integration efforts, which was crucial given the complexity of their common data model. Chen et al. chen2016agile adopted an architecture-centric approach, which strongly emphasizes designing and developing a system’s architecture as a crucial element of the development process. This approach entails the creation of a well-defined architecture, which acts as a blackprint for the system’s development. It guides the selection of technologies, allocation of resources, and risk management throughout the development life cycle. This approach complements agile’s modular design, a software development approach that emphasizes breaking down a project into smaller, more manageable modules or components to effectively manage the integration of diverse data types. The approach includes using data lakes and cloud storage to handle various data types, supporting agile’s requirement for flexible data structures and continuous delivery. The new methodology, AABA (Architecture-centric Agile Big Data Analytics), was implemented and validated through multiple case studies encompassing 10 big data analytics projects.
@@
Diverse and Decentralized Data Sources: Having diverse data sources poses a significant challenge for agile teams operating in distributed environments. Data is often fragmented across systems, teams, and geographic locations, making it difficult to consolidate and analyze effectively. This decentralization can lead to duplication of effort, delays in accessing key insights, and inconsistencies in data quality, all of which undermine the agility and responsiveness of teams. Upender upender2005staying discussed the complexity involved in integrating patient data and transitioning to a centralized data management system within a diverse and decentralized user environment. Harper and Dagnino harper2014agile highlighted the need to merge and clean data sets from different sources to enable advanced data analytics. The challenge is initiating and sustaining processes while ensuring data exchange and requests between components. The proposed solution is to distribute work requests across computing resources. Chhillar and Sharma chhillar2019act discussed the challenge of consolidating reports from various sources, such as development, testing, and bug-tracking tools, into a unified data analytics tool. This challenge pertains to seamlessly integrating diverse data sources and formats to generate real-time metrics and reports for software quality analysis.
@@ 3.2.5 Data Collection Challenges and Solutions
Table 4: Summary of Data Collection Challenges and Solutions identified by the SLR.
Challenge Solution Solution Status Ref.
@@
Capturing Diverse Data User-centered design strategies Implemented pater2018advancing
Continuous Integration tools Proposed matthies2019towards
Analytics-Driven Testing (ADT) Proposed batarseh2018predicting
@@
Data Collection Methods Automated Testing Dashboards Partially Implemented chhillar2019act
Toolchain automation Proposed lehtonen2017visualizations
Human-centred agile platform Proposed lin2018towards
Automated methods Proposed svensson2019unfulfilled
Centralized data management Proposed upender2005staying
@@
Data Sharing and Collaboration Diagnostic model Proposed fabijan2016lack
DDSE methodologies Proposed kaur2020dialogue
Include a legal advisor for compliance Proposed barbala2023data
Informative Data Collection Ontology-based approach Proposed dos2021ontology
@@
Integration project data into Retrospective Proposed matthies2020playing
Q-Rapids tool Implemented martinez2019continuously
Data analyst collaboration Proposed dharmapal2016big
@@
Comprehensive Data Collection Creation of a generalized dataset Proposed das2015towards
Big data stack Proposed huang2021leveraging
Project Management Information System Proposed fagarasan2023integrating
Automated data collection mechanisms Proposed olsson2018challenges
@@
Capturing Diverse Data: This involves collecting information from various sources and formats to meet agile development needs. In agile projects, diverse data may originate from user feedback, software metrics, and operational environments, each with unique structures and characteristics. The challenge lies in gathering this data systematically and ensuring its relevance for decision-making. Pater et al. pater2018advancing focused on capturing data for personalized healthcare and human services while addressing barriers to widespread adoption and standardization of information systems within healthcare. Matthies and Hesse matthies2019towards discussed the need for effective data collection and analysis from diverse sources, including software project data, customer feedback, and operational metrics. Furthermore, Batarseh and Gonzalez batarseh2018predicting highlighted the issue of data is often not collected in a structured and organized manner, making it difficult to analyze and use for decision-making.
@@
To address capturing diverse data challenges, Pater et al. pater2018advancing proposed a solution that involves leveraging a variety of user-centered design strategies, such as participatory design, empathic design, and design thinking, to guide the data collection methods. This approach maintains the user at the center of the process and continuously incorporates user-driven input into each phase of the agile development cycle. The solution was implemented to modernize the Army Community Service’s information technology infrastructure. Matthies and Hesse matthies2019towards highlighted tools such as Continuous Integration servers, static analysis tools, and other software development tools that provide valuable data points on the current status and health of the developed software project. They highlighted that combining data analysis and interpretation by agile teams can enable better-informed business and software development decisions. Batarseh and Gonzalez batarseh2018predicting proposed the Analytics-Driven Testing (ADT) process. ADT involves a structured data collection and analysis approach within the agile software development lifecycle. During each sprint, data are systematically collected using the ADT tool. This includes recording all software failures, measuring the Mean Time Between Failures (MTBF), and collecting other relevant data such as software module, sprint number, and run time.
@@
Data Collection Methods: Challenges related to data collection methods arise when agile teams rely on manual, error-prone, or inconsistent methods for gathering data. Agile environments demand real-time, automated, reliable data collection processes to maintain efficiency. Chhillar and Sharma chhillar2019act discussed the difficulty in generating real-time software quality metrics and the need for actual data collection processes. In contrast, Lehtonen et al. lehtonen2017visualizations discussed the manual nature of data collection from the issue management system, which can be time-consuming and susceptible to errors. Upender upender2005staying discussed the challenge involves creating a system that allows for the electronic capture of clinical research data, replacing the existing rudimentary tools such as Excel and Access that various independent research groups use. Lin et al. lin2018towards discussed the difficulty of collecting data on software engineering activities, such as task allocation, collaboration, and mood stability, without disrupting the normal Agile Software Development (ASD) process. Practitioners surveyed by Svensson et al. svensson2019unfulfilled expressed concerns regarding data availability, excessive quantity, and ambiguity regarding usability, relevance, and its connection to decision-making.